In young's double silt experiment the distance between the silt is 1.2mm and the screen is 0.75m from the silts. If the distance of the 5th fringe from the central fringe on the screen is 1.8mm. Calculate the wavelength of light used. What will be the distance of the 5th fringe from the center of the screen? 
 

Grade:12

2 Answers

Arun
4 years ago

As we learnt in

Young Double Slit Experiment -

 

- wherein

y=\Delta x\cdot \left ( \frac{D}{d} \right )

y= Distance of a point on screen from central maxima

\Delta x= Path difference at that point

 

 

Ref y be the distance from the central maximum to the point where the bright fringes due to both the wavelength coincides

For \lambda_{1},  y=\frac{m\lambda_{1}D}{d}

for  \lambda_{2},   y=\frac{n\lambda_{2}D}{d}

\therefore\ \; m \lambda_{1}=n\lambda_{2}

Vikas TU
4 years ago
Dear student 
For more info plese refer the link 
Good Luck